OpenStudy (anonymous):

A cylindrical water glass has a radius of 11 centimeters and a height of 10 centimeters. What is its volume? 3,801.33 cm3 1,267.11 cm3 5,575.28 cm3 3,455.75 cm3

OpenStudy (anonymous):

V = πr^2h

OpenStudy (anonymous):

@Calcmathlete you are on fire today

OpenStudy (anonymous):

lol. I'm just bored. Summer vacation and I have nothing to do right now. :)

OpenStudy (anonymous):

Can you show me how to do it?

OpenStudy (anonymous):

lucky my summer vacation isn't till 3 weeks from now

OpenStudy (cwtan):

Volume= Area of base x Height = \[\pi r^2 h\] Area of base = \[\pi r^2\]

OpenStudy (anonymous):

3,455.75 cm3?

OpenStudy (anonymous):

V = πr^2h V = π(11)^2(10) V = 121(10)(π) V = 1210(3.14)

OpenStudy (anonymous):

I got 3,799.4?

OpenStudy (anonymous):

Yes. I got that.

OpenStudy (anonymous):

But that's not one of the answer choicess...... @calcmathlete